OpenStudy (anonymous):

Solve the system and explain what the results mean according to your scenario. 2x + y = 11 3x + 4y = 24

OpenStudy (anonymous):

multiply the first equation by 4 and then subtract the two equation , solve for x, then substitute in anyone of them to get y!

OpenStudy (dumbcow):

Multiply top equation by -4 to eliminate y variable -8x-4y = -44 3x +4y = 24 -------------- -5x = -20 x = 4 Substitute x=4 back into original top equation 2(4) +y = 11 8+y = 11 y = 3 This means the 2 lines intersect at point (4,3) the solution, x=4, y=3 is the only x,y combination that satisfies both equations